QUESTION 2 Solve sinx = 13 2 on the interval [0, 1] 0A x = only. B.X=. 5T 6 Oc. = 11 3 D. 2.x= 2T or x = 3 2010
Answer #1

sinx=d3 x= sin ( 93 ) X122) X = n +(In T 3 120 l X = T Carl21 3 & X - T Corl x=21 3 3 option